# Exploiting Domain-Specific Parallel Data on Multilingual Language Models for Low-resource Language Translation

### v1

@@ 5.2.1 In-Domain Test Case
0cm Language Intermediate task - PMI/Gvt Final task - Bible Intermediate task - Bible Final task - PMI/Gvt Intermediate task - CC Final task - PMI/Gvt Intermediate task - CC Final task - Bible Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline si 18.9 19.4 19.4 17.0 22.1 19.8 24.5 20.2 24.1 23.7 25.8 20.2 19.3 17.4 18.8 17.0 ta 13.4 13.5 14.5 9.3 18.3 18.3 19.7 18.6 20.0 20.0 21.0 18.6 14.5 14.9 16.4 9.3 gu 14.3 14.6 15.1 12.9 19.9 20.5 21.5 20.9 23.8 24.3 24.6 20.9 13.9 13.3 15.4 12.9 hi 15.3 15.9 16.7 14.0 18.7 18.1 19.2 18.2 20.0 20.8 21.1 18.2 16.1 16.2 16.3 14.0 ka 12.7 12.8 13.8 8.2 12.9 12.5 15.5 6.5 12.9 13.8 14.8 6.5 9.6 9.6 10.0 8.2 Avg 14.9 15.2 15.9 12.3 18.4 17.8 20.1 16.9 20.2 20.5 21.5 16.9 14.7 14.3 15.4 12.3
@@
0cm Language Intermediate task - PMI/Gvt Final task - Bible Intermediate task - Bible Final task - PMI/Gvt Intermediate task - CC Final task - PMI/Gvt Intermediate task - CC Final task - Bible Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline si 38.1 37.9 37.9 37.9 44.2 43.7 44.0 44.7 44.4 43.9 44.7 44.7 37.9 38.0 37.9 37.9 ta 30.9 30.4 30.7 31.0 37.2 37.0 38.2 36.8 37.2 37.7 37.9 36.8 30.8 30.6 30.7 31.0 gu 27.4 27.7 28.2 27.8 37.6 37.7 37.6 37.9 38.3 37.5 37.6 37.9 26.9 27.8 27.9 27.8 hi 31.6 31.4 31.4 31.4 34.4 35.1 35.1 35.2 35.5 35.4 35.3 35.2 31.5 31.4 31.1 31.4 ka 27.5 26.2 27.2 27.8 32.8 32.7 34.4 33.6 34.3 33.5 33.7 33.6 27.3 26.5 27.4 27.8 Avg 31.1 30.7 31.1 31.2 37.2 37.2 37.9 37.6 37.9 37.6 37.8 37.6 30.9 30.9 31.0 31.2
@@
According to Table 5, multi-domain ITTL is the best-performing technique. However, on average, its gains over the second-best technique (multi-domain FT, which uses less compute) is less than 1 spBLEU. Hence, under limited computing resources, multi-domain FT the best option.
@@
Comparing Table 6 against Table 5 (small-small setup), increasing the intermediate task size from 1k to 25k is consistently beneficial. In this data setup, multi-domain ITTL performs better than the other two techniques (more than 1 average spBLEU gains except in one experiment) irrespective of the domain differences, thus confirming Adelani et al.’s [8] observations (see Section 2.1).
@@
The above observations are supported by Figure 2 as well 555Graphs for all the experiments are in D. Note that the line corresponding to 0k in the graphs is the vanilla FT baseline. When the target (final) task size is less than 10k, all three techniques are effective over the vanilla FT baseline, but to varying degrees. However, as the target task data set size increases, their results converge to baseline. Even using 100k CC data was not able to make gains in performance when the target domain data set is at least 25k. Thus our results emphasize that Adelani et al.’s [8] observations do not always hold, and it is important to consider the size of the datasets.
An important observation is the results for ka, which is unseen in mBART. Adding an intermediate task with 1k auxiliary data only results in a less than 1 spBLEU gain for this language in the small-small data setup. However, when adding 25k auxiliary data at the intermediate stage results in a significant gain (5.9 spBLEU on average) resulting in the largest gain amongst the languages.
5.2.2 Out-domain Test Case
@@
0mm Language Intermediate task - PMI/Gvt Final task - Bible Intermediate task - Bible Final task - PMI/Gvt Intermediate task - CC Final task - PMI/Gvt Intermediate task - CC Final task - Bible Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline Single ITTL Multi FT Multi ITTL Baseline si 7.9 11.3 11.3 0.9 5.8 4.2 5.6 3.9 8.9 9.7 10.6 3.9 3.5 8.8 8.4 0.9 ta 5.4 9.4 8.7 0.7 4.9 3.8 4.3 2.6 10.1 12.2 12.4 2.6 7.2 10.8 10.6 0.7 gu 15.3 19.8 19.7 2.2 10.5 9.3 10.2 7.3 16.8 16.1 16.5 7.3 9.6 12.2 11.9 2.2 hi 12.8 17.1 17.1 2.6 8.8 7.1 8.2 7.1 13.3 16.3 16.0 7.1 7.1 14.9 14.9 2.6 ka 7.6 14.3 13.7 0.3 4.5 4.4 5.1 1.0 4.5 4.7 5.6 1.0 0.9 1.7 1.9 0.3 Avg 9.8 14.4 14.1 1.34 6.9 5.8 6.7 4.4 10.7 11.8 12.2 4.4 5.7 9.7 9.5 1.3
@@
We use the results for the FLORES test set to explain our observations in Tables 9-12. In ITTL experiments, either PMI/Gvt or Bible is used as the final stage of fine-tuning. PMI/Gvt, Bible, and/or CC are used at the intermediate stage666Unlike domain-specific datasets s.a Bible, CC corpus is larger even for LRLs. Thus we do not run 1k setup for CC.. Here, baseline refers to vanilla FT with the dataset used for the final task (e.g. if Bible is used as the final task in the ITTL experiments, baseline refers to vanilla FT with Bible and testing with FLORES).
@@
The impact of domain divergence is visible in Figure 3 as well. For example, in single-domain ITTL in Figure 3(b), results for FLORES decrease to baseline when Bible dataset size increases in the final stage. In contrast, in Figure 3(a), when PMI/Gvt is used in the final stage, results for FLORES mostly increase when PMI/Gvt size increases.
For all four data size setups, all three techniques are better than the vanilla FT baseline, but in varying degrees. In order to further explain this observation, we calculated the correlation () between JSD (of the final stage dataset and test set) and the corresponding spBLEU score. We see that the average is for the baseline, while the same is , 0 and for single-domain ITTL, multi-domain ITTL and multi-domain FT (respectively). This indicates that, compared to the vanilla FT baseline, domain divergence is less correlated to the performance when auxiliary domain data is used.
@@
6 Ablation Study
6.1 Impact of Up-sampling
In both small-large and large-small dataset setups, one dataset size is significantly smaller than the other. In order to dilute the impact of size difference, we up-sampled777random oversampling (https://t.ly/MLrhd), we use the term up-sampling to mean oversampling the smaller data set to match with the larger dataset. For this experiment, we used CC (the largest dataset) and Bible, and experimented with three languages. As shown in Table 13, we do not see any gains from upsampling.
@@ 6.1 Impact of Up-sampling
6.2 Mixing more than two domains
@@
7 Conclusions
@@
The optimal way of using auxiliary domain parallel data to build domain-specific LRL-NMT systems on msLMs was an under-explored problem. We presented the first comparative study of two ways to utilize this parallel data—fine-tuning and additional pre-training. Based on our experiments, we are able to provide a set of observations, which will assist NMT researchers in selecting the optimal way to use data from auxiliary domains, without having to run extensive sets of experiments. In the future, we plan to experiment with multilingual FT and ITTL, as well as multilingual and multi-domain FT and ITTL. It would also be worthwhile to explore new pre-training objectives on parallel data as well.
@@ Appendix A Experiment Setup
A.1 Corpus Details (from [10])
PMIndia corpus (PMI)
[29] is a multi-way parallel corpus of 14 languages (13 Indian languages, plus English). The data set contains Indian Prime Minister’s speeches extracted from the Prime Minister of India website888https://www.pmindia.gov.in/en/.
Government corpus (Gvt)
[30] is a multi-way parallel corpus for Sinhala, Tamil, and English. This is a manually curated data set. It consists of data from official government documents (annual reports, committee reports, government institutional websites, and acts of the Parliament).
Bible corpus
We scrape Bible data from web999Sinhala: https://www.wordproject.org/bibles/si/index.htm; and others: https://ebible.org/download.php and then automatically align the sentences (at verse level) using the scripts provided by Nayak et al. [10] This curated multi-way parallel corpus has 25k parallel sentences in kn, gu, hi, ta. Note that Sinhala was scraped from a different website, thus has different content101010 We will be releasing the scripts to create the corpus on acceptance of the paper..
CCAligned (CC)
is an automatically sentence-aligned (bi-text mined) corpus based on the CommonCrawl.
FLORES-101
A.2 Model Training
Our implementation is based on the Fairseq library. All experiments are run with seed 222. We used an Nvidia Volta of 32 GPU RAM.
Training details are as follows: number of epochs =, learning rate = , dropout = , attention dropout = , and a batch size = for training and evaluation.
A.3 Evaluation Metric
spBLEU calculates the BLEU score for the text tokenized using the sentence-piece sub-word model, which has been trained for all the 101 languages in Flores-101 data set (including the languages (except kn) considered in our experiments). Furthermore, Goyal et al. [26] showed that spBLEU functions similarly to BLEU and has a strong correlation with the Chrf++ metric. [32] 121212We use the official implementation provided in the sacreBLEU library111111https://github.com/mjpost/sacreBLEU [33].

### v2

@@ 5.2.1. In-Domain Test Case
| si | 18.9 | 19.4 | 19.4 | 17.0 | 22.1 | 19.8 | 24.5 | 20.2 | 24.1 | 23.7 | 25.8 | 20.2 | 19.3 | 17.4 | 18.8 | 17.0 |
| ta | 13.4 | 13.5 | 14.5 | 9.3 | 18.3 | 18.3 | 19.7 | 18.6 | 20.0 | 20.0 | 21.0 | 18.6 | 14.5 | 14.9 | 16.4 | 9.3 |
| gu | 14.3 | 14.6 | 15.1 | 12.9 | 19.9 | 20.5 | 21.5 | 20.9 | 23.8 | 24.3 | 24.6 | 20.9 | 13.9 | 13.3 | 15.4 | 12.9 |
| hi | 15.3 | 15.9 | 16.7 | 14.0 | 18.7 | 18.1 | 19.2 | 18.2 | 20.0 | 20.8 | 21.1 | 18.2 | 16.1 | 16.2 | 16.3 | 14.0 |
| ka | 12.7 | 12.8 | 13.8 | 8.2 | 12.9 | 12.5 | 15.5 | 6.5 | 12.9 | 13.8 | 14.8 | 6.5 | 9.6 | 9.6 | 10.0 | 8.2 |
| Avg | 14.9 | 15.2 | 15.9 | 12.3 | 18.4 | 17.8 | 20.1 | 16.9 | 20.2 | 20.5 | 21.5 | 16.9 | 14.7 | 14.3 | 15.4 | 12.3 |
@@
| si | 38.1 | 37.9 | 37.9 | 37.9 | 44.2 | 43.7 | 44.0 | 44.7 | 44.4 | 43.9 | 44.7 | 44.7 | 37.9 | 38.0 | 37.9 | 37.9 |
| ta | 30.9 | 30.4 | 30.7 | 31.0 | 37.2 | 37.0 | 38.2 | 36.8 | 37.2 | 37.7 | 37.9 | 36.8 | 30.8 | 30.6 | 30.7 | 31.0 |
| gu | 27.4 | 27.7 | 28.2 | 27.8 | 37.6 | 37.7 | 37.6 | 37.9 | 38.3 | 37.5 | 37.6 | 37.9 | 26.9 | 27.8 | 27.9 | 27.8 |
| hi | 31.6 | 31.4 | 31.4 | 31.4 | 34.4 | 35.1 | 35.1 | 35.2 | 35.5 | 35.4 | 35.3 | 35.2 | 31.5 | 31.4 | 31.1 | 31.4 |
| ka | 27.5 | 26.2 | 27.2 | 27.8 | 32.8 | 32.7 | 34.4 | 33.6 | 34.3 | 33.5 | 33.7 | 33.6 | 27.3 | 26.5 | 27.4 | 27.8 |
| Avg | 31.1 | 30.7 | 31.1 | 31.2 | 37.2 | 37.2 | 37.9 | 37.6 | 37.9 | 37.6 | 37.8 | 37.6 | 30.9 | 30.9 | 31.0 | 31.2 |
@@
According to Table 5, multi-domain ITTL is the best-performing technique. However, on average, its gains over the second-best technique (multi-domain FT, which uses less compute) is less than 1 spBLEU. Hence, under limited computing resources, multi-domain FT is the best option.
@@
Comparing Table 6 against Table 5 (small-small setup), increasing the intermediate task size from 1k to 25k is consistently beneficial. In this data setup, multi-domain ITTL performs better than the other two techniques irrespective of the domain differences, thus confirming Adelani et al.’s (Adelani et al., 2022) observations (see Section 2.1).
@@
The above observations are supported by Figure 2 as well 666Graphs for all the experiments are in D.. Note that the line corresponding to 0k in the graphs is the vanilla FT baseline. When the target (final) task size is less than 10k, all three techniques are effective over the vanilla FT baseline, but to varying degrees. However, as the target task data set size increases, their results converge to the baseline. Even using 100k CC data was not able to make gains in performance when the target domain data set is at least 25k. Thus our results emphasize that Adelani et al.’s (Adelani et al., 2022) observations do not always hold, and it is important to consider the size of the datasets.
An important observation is the results for ka, which is unseen in mBART. Adding an intermediate task with 1k auxiliary data only results in a less than 1 spBLEU gain for this language in the small-small data setup. However, when adding 25k auxiliary data at the intermediate stage results in a significant gain (5.9 spBLEU on average) resulting in the largest gain among the languages.
In summary, for the small-small setup, multi-domain FT is the best option. For the large-small setup, multi-domain ITTL is the best. In both the small-large setup and the large-large setup, vanilla FT baseline is the best. Contrary to what Adelani et al. (Adelani et al., 2022) reported, our experiments confirm that the best model depends on the dataset size.
5.2.2. Out-domain Test Case
@@
| si | 7.9 | 11.3 | 11.3 | 0.9 | 5.8 | 4.2 | 5.6 | 3.9 | 8.9 | 9.7 | 10.6 | 3.9 | 3.5 | 8.8 | 8.4 | 0.9 |
| ta | 5.4 | 9.4 | 8.7 | 0.7 | 4.9 | 3.8 | 4.3 | 2.6 | 10.1 | 12.2 | 12.4 | 2.6 | 7.2 | 10.8 | 10.6 | 0.7 |
| gu | 15.3 | 19.8 | 19.7 | 2.2 | 10.5 | 9.3 | 10.2 | 7.3 | 16.8 | 16.1 | 16.5 | 7.3 | 9.6 | 12.2 | 11.9 | 2.2 |
| hi | 12.8 | 17.1 | 17.1 | 2.6 | 8.8 | 7.1 | 8.2 | 7.1 | 13.3 | 16.3 | 16.0 | 7.1 | 7.1 | 14.9 | 14.9 | 2.6 |
| ka | 7.6 | 14.3 | 13.7 | 0.3 | 4.5 | 4.4 | 5.1 | 1.0 | 4.5 | 4.7 | 5.6 | 1.0 | 0.9 | 1.7 | 1.9 | 0.3 |
| Avg | 9.8 | 14.4 | 14.1 | 1.34 | 6.9 | 5.8 | 6.7 | 4.4 | 10.7 | 11.8 | 12.2 | 4.4 | 5.7 | 9.7 | 9.5 | 1.3 |
@@
We use the results for the FLORES test set to explain our observations in Tables 9-12. In ITTL experiments, either PMI/Gvt or Bible is used as the final stage of fine-tuning. PMI/Gvt, Bible, and/or CC are used at the intermediate stage777Unlike domain-specific datasets s.a Bible, CC corpus is larger even for LRLs. Thus we do not run 1k setup for CC.. Here, baseline refers to vanilla FT with the dataset used for the final task (e.g. if Bible is used as the final task in the ITTL experiments, baseline refers to vanilla FT with Bible and testing with FLORES).
@@
The impact of domain divergence is visible in Figure 3 as well. For example, in Figure 3(a), when PMI/Gvt is used in the final stage, results for FLORES mostly increase when PMI/Gvt size increases. In contrast, in the single-domain ITTL in Figure 3(b), results for FLORES decrease to baseline when Bible dataset size increases in the final stage.
For all four data size setups, all three techniques are better than the vanilla FT baseline, but in varying degrees. In order to further explain this observation, we calculated the correlation () between JSD (of the final stage dataset and test set) and the corresponding spBLEU score. For the baseline, the average is , while the same is , 0 and for single-domain ITTL, multi-domain ITTL and multi-domain FT (respectively). This indicates that, compared to the vanilla FT baseline, domain divergence is less correlated to the performance when auxiliary domain data is used.
@@
In summary, all three techniques (single ITTL, multi-FT and multi-ITTL) outperform the vanilla FT baseline for all the data sizes. Among the three strategies, multi-FT and multi-ITTL outperform single ITTL in general. However, there is no clear winner between multi-FT and multi-ITTL. Considering the less computational cost involved with multi-FT and its less susceptibility to domain divergence, multi-FT would be the better choice than multi-ITTL. However, the exact results of all the strategies depend on the divergence between domains. Therefore, domains closer to the target domain should be selected to train the msLM.
6. Ablation Study
6.1. Impact of Up-sampling
In both small-large and large-small dataset setups, one dataset size is significantly smaller than the other. In order to dilute the impact of size difference, we up-sampled888random oversampling (https://t.ly/MLrhd), we use the term up-sampling to mean oversampling the smaller data set to match with the larger dataset. For this experiment, we used CC (the largest dataset) and Bible, and experimented with three languages. As shown in Table 13, we do not see any gains from upsampling.
@@ 6.1. Impact of Up-sampling
6.2. Mixing more than two domains
@@
| si | 12.00 | 12.25 | 12.04 | 12.08 | 24.18 | 23.99 | 24.19 | 24.15 |
| ta | 8.65 | 8.87 | 8.44 | 8.75 | 14.55 | 14.91 | 14.75 | 14.96 |
| gu | 10.84 | 11.34 | 11.09 | 11.09 | 16.91 | 18.08 | 16.94 | 17.36 |
| hi | 18.53 | 19.03 | 18.41 | 18.49 | 24.49 | 25.49 | 24.15 | 24.99 |
| ka | 6.27 | 6.80 | 6.47 | 6.42 | 14.85 | 15.07 | 14.91 | 15.45 |
| Avg | 11.26 | 11.66 | 11.29 | 11.37 | 19.00 | 19.51 | 18.99 | 19.38 |
| si | 5.88 | 8.63 | 7.68 | 5.26 | 7.13 | 8.63 | 7.67 | 7.94 |
| ta | 5.41 | 8.47 | 7.00 | 6.49 | 7.13 | 8.47 | 7.70 | 7.78 |
| gu | 13.12 | 15.24 | 14.86 | 13.38 | 13.74 | 15.24 | 14.40 | 15.11 |
| hi | 17.02 | 21.55 | 21.08 | 18.86 | 19.62 | 21.55 | 20.12 | 21.38 |
| ka | 6.91 | 11.10 | 9.31 | 7.05 | 9.17 | 11.10 | 9.64 | 10.44 |
| Avg | 9.67 | 13.00 | 11.99 | 10.21 | 11.36 | 13.00 | 11.91 | 12.53 |
6.3. Impact of the msLM
In order to determine whether our observations hold across msLMs, we conducted an ablation study with the NLLB-200-distilled-600M model for both the in-domain and the out-domain cases, considering the small-small data setup. Results are reported in Tables 16 and 17. Here, Table 16 has the same setup as Table 5 for the in-domain case, and Table 17 has the same setup as Table 9 for the out-domain case. In this data setup, reaffirming the recommendation we made for mBART, multiFT emerges as the most promising technique for the in-domain case, despite there being a lesser results variation among the techniques compared to that of mBART. As for the out-domain case, the observations made with the mBART models are mostly echoed with NLLB as well. To begin with, vanilla FT baseline performs better when PMI/Gvt is used for fine-tuning, compared to using Bible. Similarly, the best results from Single ITTL are reported when a domain closer to FLORES is used in the final fine-tuning stage. However, for Multi ITTL, there is no result variation with respect to the domain used in the final stage. Overall, both multi-domain ITTL and multi-domain FT are better than single-domain ITTL. However, some individual ITTL results fall below the baseline. Here also we can conclude that multiFT is better than multiITTL.
7. Discussion
7.1. Recommendations
Based on our extensive experiments, the following recommendations can be made:
When the parallel data set size is small (less than 25k), do not pre-train mBART with bitext denoising, as it may not result in any gains.
In the in-domain case, when two domains are concerned, and (1) if both have small amounts of data (about 1k), use the multi-domain FT option, (2) use the multi-domain ITTL option when the intermediate dataset size increases, (3) Simply use vanilla FT if the target domain has at least 25k data samples.
In the out-domain test case, when two domains are concerned, results are heavily impacted by the divergence of domains. Therefore, before carrying any experiments, calculate the divergence between domains and use data from a domain that is closest to the target domain. We recommend multi-FT in this case, since its results are less susceptible to domain divergence and it is computationally more efficient.
Do not blindly mix data belonging to more than two domains, as the effective gains are highly subjective to the domain divergence.
7.2. Limitations
In our experiments, we considered only mBART, with an ablation study on NLLB. Other msLMs could not be considered due to the limitations in computational resources. Similarly, the number of languages had to be kept at 5, due to the difficulty in finding data belonging to multiple domains. Due to this, it was not possible to carry out a language-wise analysis. The maximum dataset size for most domains/languages was 25k, a common problem with LRLs. To keep the results discussion concise, only a selected set of EN-XX experiment results were used for the discussion. For clarity, only spBLEU results are reported. We could not report results on human evaluations, due to the lack of funds to employ professional translators.
8. Conclusion
@@
The optimal way of using auxiliary domain parallel data to build domain-specific LRL-NMT systems on msLMs was an under-explored problem. We presented the first comparative study of two ways to utilize this parallel data—fine-tuning and additional pre-training. Our key findings are: (1) with small parallel data ( 25k per domain in our experiments), continuous pre-training with bitext denoising does not provide visible gains; (2) for in-domain translation, multi-domain FT is better for small amounts of data and multi-domain ITTL is the best when the dataset size increases, however when the target domain has 25k parallel sentences, vanilla FT is usually sufficient; (3) for out-of-domain evaluation, auxiliary-domain strategies generally improve over vanilla FT, but the best choice depends on the divergence between auxiliary and target domains; (4) multi-domain FT offers a strong performance/compute trade-off and is generally less sensitive to domain divergence than multi-domain ITTL; and (5) blindly mixing data from multiple domains do not deliver expected results due to domain divergence. We hope our findings and recommendation will assist NMT researchers in selecting the optimal way to use data from auxiliary domains, without having to run extensive sets of experiments. In the future, we plan to experiment with multilingual FT and ITTL, as well as multilingual and multi-domain FT and ITTL. It would also be worthwhile to explore new pre-training objectives on parallel data as well.
@@ Appendix A Experiment Setup
A.1. Corpus Details (from (Nayak et al., 2023))
PMIndia corpus (PMI) (Haddow and Kirefu, 2020) is a multi-way parallel corpus of 14 languages (13 Indian languages, plus English). The data set contains Indian Prime Minister’s speeches extracted from the Prime Minister of India website999https://www.pmindia.gov.in/en/.
Government corpus (Gvt) (Fernando et al., 2020) is a multi-way parallel corpus for Sinhala, Tamil, and English. This is a manually curated data set. It consists of data from official government documents (annual reports, committee reports, government institutional websites, and acts of the Parliament).
Bible corpus We scrape Bible data from web101010Sinhala: https://www.wordproject.org/bibles/si/index.htm; and others: https://ebible.org/download.php and then automatically align the sentences (at verse level) using the scripts provided by Nayak et al. (Nayak et al., 2023) This curated multi-way parallel corpus has 25k parallel sentences in kn, gu, hi, ta. Note that Sinhala was scraped from a different website, thus has different content111111We will be releasing the scripts to create the corpus on acceptance of the paper..
CCAligned (CC) is an automatically sentence-aligned (bi-text mined) corpus based on the CommonCrawl.
FLORES-101 (Goyal et al., 2022) is a multi-way parallel corpus. The source is English Wikipedia, which has been manually translated into 101 languages. Data comes from a variety of topics and domains, thus can be considered open-domain. We use Floresv1 (Guzmán et al., 2019) for Sinhala since it is not present in Flores-101.
When the size of a training dataset exceeded the amount of samples needed for the experiment, we randomly sampled data from the source corpus.
A.2. Model Training
The mBART experiments were implemented using the Fairseq library and conducted on an NVIDIA Volta GPU with 32 GB of memory.
The NLLB experiments were implemented using the Hugging Face Transformers library and conducted on an NVIDIA H100 GPU with 80 GB of memory.
All experiments were conducted with a fixed random seed of 222.
Training details are as follows: number of epochs =, learning rate = , dropout = , attention dropout = , and a batch size = for training and evaluation. mBART tokenizer was used to tokenize text.
A.3. Evaluation Metric
spBLEU calculates the BLEU score for the text tokenized using the sentence-piece sub-word model, which has been trained for all the 101 languages in Flores-101 data set (including the languages (except kn) considered in our experiments). Furthermore, Goyal et al. (Goyal et al., 2022) showed that spBLEU functions similarly to BLEU and has a strong correlation with the Chrf++ metric. (Popović, 2017) 131313We use the official implementation provided in the sacreBLEU library121212https://github.com/mjpost/sacreBLEU (Post, 2018).